# Are formal and functional linguistic mechanisms dissociated in language models?

### v3

@@ 6.3 What components overlap between formal and functional tasks?
Our previous experiments suggest that there is formal-functional overlap, even when considering functional tasks containing no linguistic structure. But what could this overlap consist of? In order to capture entire model mechanisms for a given task, circuits must be whole paths from models’ inputs to logits (Section 3). This means that, if there exist low-level task-agnostic mechanisms shared between formal and functional circuits, these might constitute a formal-functional overlap. For example, prior work suggests that transformer LMs’ early layers help detokenize and contextualize words (Ghandeharioun et al., 2024). These could act as a shared low-level input processing stream, after which distinct formal or functional modules take charge.
Such low-level cross-task mechanisms also exist in the brain: brain regions responsible for vision, for example, must be active whether reading sentences or non-words. However, the localizer approach used to find the language network (Section 2.1) naturally avoids this issue by looking at the difference in activations in brain regions in each condition. Reading-relevant visual brain regions should activate equally in the sentence and non-word conditions, so they are not captured by the language localizer.
@@
We find that the components and edges that are relevant across circuits are dominated by MLPs: all MLPs are included in every circuit. Our edge-level analysis (Figure 6, left) shows that the vast majority of edges shared across circuits involve MLPs. This is despite the fact that attention heads are far more numerous within all models. However, Figure 6 (right) shows that the components and edges shared between circuits, including the MLPs involved, span all layers. So, insofar as a shared low-level processing area should be located at early layers of the model only (or indeed layer-wise localized at all), this hypothesis is false. Rather, this shared network seems to be composed of an “MLP backbone”, running up and down the model, which is essential for its functioning. Unfortunately, the role of MLPs in general is rather contested; while they have been implicated in detokenization, they have also been conceived of as key-value memories for fact storage (Geva et al., 2021), and implicated in various task-specific roles (Hanna, Liu, and Variengien, 2023; Lieberum et al., 2023; Nikankin et al., 2024).
Finally, when we exclude this shared network from our overlap analysis, the size of each circuit is sharply reduced: circuits’ sizes fall to half of their previous size, or less. However, some overlap remains: the median IoU shrinks, but only from 0.11 to 0.10. Moreover, as we have excluded the same shared network from every circuit, the trends in IoU and recall, regarding which circuits overlap the most, remain the same. Overall, we can say that this shared MLP backbone is what is most drives overlap, though it does not explain all of it.
@@ 7 Are Formal and Functional Task Circuits Cross-Task Faithful?
Examining the pair-wise cross-task faithfulness of formal and functional tasks (Figure 7, left), different trends emerge compared to earlier. While Capital-Country and Country-Capital are still similar, the IOI / Colored Objects / Entity Tracking grouping is much less clear. Moreover, while cross-task faithfulness is somewhat influenced by circuit size—the tasks whose circuits best perform other tasks (i.e., Hypernymy, Colored Objects, whose columns are dark) are large—other tasks with large circuits, like Entity Tracking do not perform other tasks as well In general, the median cross-task faithfulness between two distinct formal tasks (0.11), or two distinct functional tasks (0.14) is higher than the median in formal-functional or functional-formal conditions (0.01 and 0.05), but this difference is still not great.
However, the results of our clustering analysis (Figure 7, right) provide more positive evidence. In this analysis, we cluster the cross-task similarity reference vectors, i.e. the vectors that show for a given task, which task’s circuits solve it best. Surprisingly, all formal tasks form one cluster that is separate all functional tasks: Gendered Pronoun, Wug Test, NPI, SVA, and Hypernymy all cluster separately from Greater-Than, IOI, Colored Objects, and Entity Tracking. We note that Capital-Country and Country-Capital form a two-task cluster that is separate from both formal and functional tasks, but this is compatible with formal-functional dissociation and formal-formal consistency. Functional tasks need not cluster together; they need only be separate from formal tasks. Overall, these results suggest that formal tasks are more alike one another with respect to which circuits solve them. So, while there is no consistent formal region at the IoU overlap level, formal task circuits are indeed more similar to each other than functional task circuits are to them at the cross-task faithfulness level. These are promising results, but to ensure they are robust, we repeat our circuit analyses at different levels of granularity in the following section.
@@ 8 Node and Neuron-Level Circuits
To answer this, we perform our analyses again, but at the node and neuron level. We adapt EAP-IG, which produces estimates of each edge’s indirect effect, to produce estimates of each node or neuron’s indirect effect. We do so by replacing $\nabla_{v_{in}}$ in Equation 3 with $\nabla_{u_{out}}$, yielding (scalar) node IE estimates:
@@
Our node-level IoU results (Figure 8, left) are rather similar to our edge-level results. The same circuits overlap with one another: Capital-Country and Country-Capital have highly similar circuits, while IOI, Colored Objects, and Entity Tracking are all similar to one another. All other circuits have low IoU. Overall, the circuits are much larger than in the edge scenario, including around 25% of the model on average. This is unsurprising, as nodes are much larger units than edges; including one node is essentially equivalent to including all of its outgoing edges. While we could previously only include one edge out of a node, if that edge was important, we now must include the entire node.
Neuron circuits (Figure 8, right) contrast more with edge circuits. The increased granularity of the computational graph in the neuron case does not enable greater sparsity; neuron circuits are more similar in size (in terms of percentage of the whole model) to node than edge circuits. While some of the same trends in task overlap are visible as in past IoU analyses, neuron circuits have low IoU in general. Thus, neither IoU analysis provides further evidence for the existence of a formal linguistic region.
@@
However, the cross-task faithfulness of node and neuron circuits (Figure 9) are more suggestive of a formal-functional distinction. In particular, formal tasks have a higher cross-task faithfulness on other formal tasks than on functional tasks, while functional tasks capture one another much better than they capture formal tasks. On the node level, the NPI and SVA circuits capture not only each other, but also the Wug Test and Gendered Pronoun tasks relatively well; on the neuron level, formal-formal similarity is more generalized, but still visible.
@@
A clustering analysis at either level (Figure 10) yields the same conclusion. At both the node and neuron level, there is a clustering of the formal tasks as separate from the functional ones. The consistent emergence of a formal grouping as measured by cross-task faithfulness provides stronger evidence for a division between formal and functional task mechanisms in LLMs. However, the formal grouping that we uncover is defined not by the precise edges, nodes, or neurons that fall within it, but by the general ability of formal tasks to perform other formal tasks.
@@ 9 Discussion
In this paper, we have translated Mahowald et al.’s (2024) hypothesis about emergent formal-functional dissociation in LLMs into the circuit analysis framework, and tested the hypothesis. We have tested this in two ways, first measuring if formal and functional circuits contain overlapping units, and then measuring the cross-task faithfulness between the two. Our results indicate that formal and functional circuits have low but non-zero overlap. Moreover, formal circuits also have low overlap in general with one another, suggesting that there is no undifferentiated formal language region in LLMs as in the human brain. Formal circuits are also not a subset of functional ones in general, even when these functional tasks are posed using language. These results are stable across different circuit granularities. However, when we measure cross-task faithfulness, a different pattern emerges. At the edge level, we see some clustering of formal tasks as separate from functional tasks. Then, at the node and neuron level, a stronger clustering of such tasks emerges, with formal tasks clustering entirely separately from functional ones.
@@
The ideal metric for mechanistic similarity should measure not only if the same units are involved in a given pair of tasks, but also if they are involved to the same extent. This could be achieved by using, e.g. a weighted recall metric, where one task’s recall of another’s edges is weighted by the IE that the latter assigns to each edge. Even better metrics could compute quantities such as the graph edit distance between the entire IE-scored computational graphs for two different tasks; this sort of metric would remove the need for choosing a circuit of a fixed size, and instead directly compare the causal relevance of every unit between the two tasks. However, as discussed in Section 5, such metrics are hampered by the fact that EAP and EAP-IG estimate IEs relatively poorly in absolute terms (Syed, Rager, and Conmy, 2024).
@@
In this paper, we have also conducted a circuit study at a larger scale than most circuit analyses. In particular, many recent circuit papers study one or two tasks in GPT-2 small (Li and Janson, 2024) or single models with larger parameter counts (Prakash et al., 2024; Bhaskar et al., 2024). In contrast, we study 5 different models of 2-8 billion parameters, across 10 tasks. This allows us to both build on existing results—we find that the IOI and Colored Objects tasks are not only similar to each other, as found by Merullo, Eickhoff, and Pavlick (2024), but also similar to Entity Tracking. Other results, such as the existence of an MLP backbone across circuits, and the fact that circuits have low overlap with one another in general, were possible only because of the larger models and more numerous tasks that we studied.
@@
In conclusion, we have investigated the question of formal-functional dissociation in LLMs via a wide-ranging circuit similarity study across five formal and five functional tasks, and three granularities. In doing so we provide the first application of circuits to questions from neuroscience and discover new phenomena, such as an MLP backbone running down all circuits. However, as in prior work, we find major differences in similarity depending on whether we measure overlap and cross-task faithfulness (Hanna, Pezzelle, and Belinkov, 2024). Measuring overlap shows low formal-functional similarity, but also low formal-formal similarity; in contrast, cross-task faithfulness suggests clear formal-functional dissimilarity and formal-formal similarity. Future work investigating more tasks could establish the range of tasks for which this holds, and further the use of mechanistic interpretability to gain a causally-informed low-level understanding of LLMs and their relationship with the brain.
@@ 10 Limitations
In this study, we analyzed five formal and five functional tasks, limited by the difficulty of crafting tasks that fit into the circuits framework; studying more tasks would help confirm that formal and functional tasks do indeed have distinct patterns in cross-task faithfulness. Moreover, we have studied circuits for these tasks found at a fixed faithfulness threshold, 85%; varying this threshold could cause our results to change.
@@ Acknowledgments
The authors would like to thank Kyle Mahowald and Anna Ivanova for helpful discussions about this paper, and Raquel Fernández for reading it and providing insightful feedback. The authors also thank the Dialogue Modeling Group at the University of Amsterdam, and the Technion NLP group for their many useful comments and suggestions. MH is supported in part by an OpenAI Superalignment Grant. This work was conducted in part during a research visit funded by the ELIAS Mobility Program.
@@
However, for some tasks (Gendered Pronoun, SVA, NPI, and Greater-Than), there is no one correct answer, and that answer need not always be the top output. Consider the case of SVA with the input The keys on the cabinet…; while are is clearly more correct than is, or any other singular-conjugated verbs, much of the model’s probability may be distributed to words like that. Such words are neither right nor wrong, so we should not penalize models for outputting them. As such we measure accuracy by scoring each example correct if models assign more probability to the right token (or set thereof) than to the wrong token (or set thereof); alternatively, we could use accuracy with respect to the first token that is either correct or incorrect, rather than neither.
Results for the tasks shared across all models are in Table 3. Accuracies tend high, except for IOI and Entity Tracking.

### v4

@@ 6.3 What components overlap between formal and functional tasks?
Our previous experiments suggest that there is formal-functional overlap, even when considering functional tasks containing no linguistic structure. But what could this overlap consist of? In order to capture entire model mechanisms for a given task, circuits must be whole paths from models’ inputs to logits (Section˜3). This means that, if there exist low-level task-agnostic mechanisms shared between formal and functional circuits, these might constitute a formal-functional overlap. For example, prior work suggests that transformer LMs’ early layers help detokenize and contextualize words (Ghandeharioun et al., 2024). These could act as a shared low-level input processing stream, after which distinct formal or functional modules take charge.
Such low-level cross-task mechanisms also exist in the brain: brain regions responsible for vision, for example, must be active whether reading sentences or non-words. However, the localizer approach used to find the language network (Section˜2.1) naturally avoids this issue by looking at the difference in activations in brain regions in each condition. Reading-relevant visual brain regions should activate equally in the sentence and non-word conditions, so they are not captured by the language localizer.
@@
We find, much like Bhaskar, Friedman, and Chen (2024), that there is indeed a set of the components and edges that are relevant across circuits. These are dominated by MLPs: all MLPs are included in every circuit. Our edge-level analysis (Figure˜6, left) shows that the vast majority of edges shared across circuits involve MLPs. This is despite the fact that attention heads are far more numerous within all models. However, Figure˜6 (right) shows that the components and edges shared between circuits, including the MLPs involved, span all layers. So, insofar as a shared low-level processing area should be located at early layers of the model only (or indeed layer-wise localized at all), this hypothesis is false. Rather, this shared network seems to be composed of an “MLP backbone”, running up and down the model, which is essential for its functioning. Unfortunately, the role of MLPs in general is rather contested; while they have been implicated in detokenization, they have also been conceived of as key-value memories for fact storage (Geva et al., 2021), and implicated in various task-specific roles (Hanna, Liu, and Variengien, 2023; Lieberum et al., 2023; Nikankin et al., 2024).
Finally, when we exclude this shared network from our overlap analysis, the size of each circuit is sharply reduced: circuits’ sizes fall to half of their previous size, or less. However, some overlap remains: the median IoU shrinks, but only from 0.11 to 0.10. Moreover, as we have excluded the same shared network from every circuit, the trends in IoU and recall, regarding which circuits overlap the most, remain the same. Overall, we can say that this shared MLP backbone is what most drives overlap, though it does not explain all of it.
@@ 7 Are formal and functional task circuits cross-task faithful?
Examining the pair-wise cross-task faithfulness of formal and functional tasks (Figure˜7, left), different trends emerge compared to earlier. While Capital-Country and Country-Capital are still similar, the IOI / Colored Objects / Entity Tracking grouping is much less clear. Moreover, while cross-task faithfulness is somewhat influenced by circuit size—the tasks whose circuits best perform other tasks (i.e., Hypernymy, Colored Objects, whose columns are dark) are large—other tasks with large circuits, like Entity Tracking do not perform other tasks as well. In general, the median cross-task faithfulness between two distinct formal tasks (0.11), or two distinct functional tasks (0.14) is higher than the median in formal-functional or functional-formal conditions (0.01 and 0.05), but this difference is still not great.
However, the results of our clustering analysis (Figure˜7, right) provide more positive evidence. In this analysis, we cluster the cross-task similarity reference vectors, i.e. the vectors that show for a given task, which task’s circuits solve it best. Surprisingly, all formal tasks form one cluster that is separate from all functional tasks: Gendered Pronoun, Wug Test, NPI, SVA, and Hypernymy all cluster separately from Greater-Than, IOI, Colored Objects, and Entity Tracking. We note that Capital-Country and Country-Capital form a two-task cluster that is separate from both formal and functional tasks, but this is compatible with formal-functional dissociation and formal-formal consistency. Functional tasks need not cluster together; they need only be separate from formal tasks. Overall, these results suggest that formal tasks are more alike one another with respect to which circuits solve them. So, while there is no consistent formal region at the IoU overlap level, formal task circuits are indeed more similar to each other than functional task circuits are to them at the cross-task faithfulness level. These are promising results, but to ensure they are robust, we repeat our circuit analyses at different levels of granularity in the following section.
@@ 8 Node and Neuron-Level Circuits
To answer this, we perform our analyses again, but at the node and neuron level. We adapt EAP-IG, which produces estimates of each edge’s indirect effect, to produce estimates of each node or neuron’s indirect effect. We do so by replacing $\nabla_{v_{in}}$ in Equation˜3 with $\nabla_{u_{out}}$, yielding (scalar) node IE estimates:
@@
Note that a neuron here refers to an individual dimension of token embedding, or an individual dimension of an attention head’s or MLP’s outputs, prior to the addition of the residual connection. This is in contrast to, e.g., a hidden neuron within an MLP.
@@
Our node-level IoU results (Figure˜8, left) are rather similar to our edge-level results. The same circuits overlap with one another: Capital-Country and Country-Capital have highly similar circuits, while IOI, Colored Objects, and Entity Tracking are all similar to one another. All other circuits have low IoU. Overall, the circuits are much larger than in the edge scenario, including around 25% of the model on average. This is unsurprising, as nodes are much larger units than edges; including one node is essentially equivalent to including all of its outgoing edges. While we could previously only include one edge out of a node, if that edge was important, we now must include the entire node.
Neuron circuits (Figure˜8, right) contrast more with edge circuits. The increased granularity of the computational graph in the neuron case does not enable greater sparsity; neuron circuits are more similar in size (in terms of percentage of the whole model) to node than edge circuits. While some of the same trends in task overlap are visible as in past IoU analyses, neuron circuits have low IoU in general. Thus, neither IoU analysis provides further evidence for the existence of a formal linguistic region.
@@
However, the cross-task faithfulness of node and neuron circuits (Figure˜9) are more suggestive of a formal-functional distinction. In particular, formal tasks have a higher cross-task faithfulness on other formal tasks than on functional tasks, while functional tasks capture one another much better than they capture formal tasks. On the node level, the NPI and SVA circuits capture not only each other, but also the Wug Test and Gendered Pronoun tasks relatively well; on the neuron level, formal-formal similarity is more generalized, but still visible.
@@
A clustering analysis at either level (Figure˜10) yields the same conclusion. At both the node and neuron level, there is a clustering of the formal tasks as separate from the functional ones. The consistent emergence of a formal grouping as measured by cross-task faithfulness provides stronger evidence for a division between formal and functional task mechanisms in LLMs. However, the formal grouping that we uncover is defined not by the precise edges, nodes, or neurons that fall within it, but by the general ability of formal tasks to perform other formal tasks.
@@ 9 Discussion
In this paper, we have translated Mahowald et al.’s (2024) hypothesis about emergent formal-functional dissociation in LLMs into the circuit analysis framework, and tested the hypothesis. We have tested this in two ways, first measuring if formal and functional circuits contain overlapping units, and then measuring the cross-task faithfulness between the two. Our results indicate that formal and functional circuits have low but non-zero overlap. Moreover, formal circuits also have low overlap in general with one another, suggesting that there is no undifferentiated formal language region in LLMs as in the human brain. Formal circuits are also not a subset of functional ones in general, even when these functional tasks are posed using language. These results are stable across different circuit granularities and faithfulness thresholds. However, when we measure cross-task faithfulness, a different pattern emerges. At the edge level, we see some clustering of formal tasks as separate from functional tasks. Then, at the node and neuron level, a stronger clustering of such tasks emerges, with formal tasks clustering entirely separately from functional ones.
@@
The ideal metric for mechanistic similarity should measure not only if the same units are involved in a given pair of tasks, but also if they are involved to the same extent. This could be achieved by using, e.g. a weighted recall metric, where one task’s recall of another’s edges is weighted by the IE that the latter assigns to each edge. Even better metrics could compute quantities such as the graph edit distance between the entire IE-scored computational graphs for two different tasks; this sort of metric would remove the need for choosing a circuit of a fixed size, and instead directly compare the causal relevance of every unit between the two tasks. However, as discussed in Section˜5, such metrics are hampered by the fact that EAP and EAP-IG estimate IEs relatively poorly in absolute terms (Syed, Rager, and Conmy, 2024).
@@
Still, even if formal-functional groupings emerge when measuring cross-task faithfulness, should this then be considered evidence for separate formal and functional modules? To answer this, we can again return to the human neuroscience literature, in which a similar phenomenon has been observed. The human brain, too, can exhibit degeneracy, whereby multiple neural circuits lead to the same downstream behavior (Price and Friston, 2002; Mason, 2010; Mason et al., 2015); this is also discussed in philosophy as multiple realizability (Bickle, 2020). Crucially, this degeneracy can be taken as evidence against modularity; if a neural module is an area of the brain that is necessary and sufficient to complete a task, then the existence of multiple circuits that are sufficient (but not on their own necessary) to complete a task provides evidence against modularity (Zerilli, 2019).
Past LLM circuit work has already shown evidence for degeneracy via the phenomenon of self-repair, where model components that initially appeared unimportant to a task take on new roles and compensate for the task’s circuit, when it is ablated (Wang et al., 2023; McGrath et al., 2023). If we interpret cross-task faithfulness as degeneracy, this serves as evidence against the existence of separate modules for each formal task. However, there does seem to be a collection of model components that support formal linguistic competence. Not all of these components are strictly necessary for each task; sometimes a task can be solved via multiple combinations of such components. This runs contrary to the idea of a formal linguistic network that is undifferentiated and wholely necessary for all formal linguistic competences. At the same time, we found that circuits for formal linguistic tasks are better at solving other formal linguistic tasks than functional linguistic circuits are.
Finally, it is worth noting that, throughout this study, we have operated under the assumption that the formal-functional distinction in the brain as discussed by Mahowald et al. is clear-cut. However, this is not uncontroversial: both the specific aspects of language that are included in the formal language network, as well as the formal-functional distinction more broadly, are both topics of heated debate. The distinction between lexical / combinatorial semantics and world knowledge is one site of conflict, where what constitutes each group, and if these groups should be separate at all, is contested (Hagoort et al., 2004; Pylkkänen, Oliveri, and and, 2009). Others reject the formal-functional distinction wholesale due to its reliance on fMRI studies, as opposed to more temporally sensitive techniques (Murphy and Woolnough, 2024), and advocate for an approach to language in the brain that stresses the connectivity between core language regions and other cognitive systems (Forkel and Hagoort, 2024). While the critiques of hypernymy’s categorization find some support in our edge IoU analyses, arguments against the formal-functional distinction more broadly are less supported by our analyses, which show some dissociation between the two.
@@
Beyond our results on formal-functional dissociation, this present study also presents a much larger-scale work than most circuit analyses. In particular, recent circuit papers often study one or two tasks in GPT-2 small (Li and Janson, 2024) or single models with larger parameter counts (Prakash et al., 2024; Bhaskar et al., 2024). In contrast, we study 5 different models of 2-8 billion parameters, across 10 tasks. This allows us to both build on existing results—we find that the IOI and Colored Objects tasks are not only similar to each other, as found by Merullo, Eickhoff, and Pavlick (2024), but also similar to Entity Tracking. Other results, such as the existence of an MLP backbone across circuits, and the fact that circuits have low overlap with one another in general, were possible only because of the larger models and more numerous tasks that we studied.
@@
In conclusion, we have investigated the question of formal-functional dissociation in LLMs via a wide-ranging circuit similarity study across five formal and five functional tasks, and three granularities. In doing so we provide the first application of circuits to questions from neuroscience and discover new phenomena, such as an MLP backbone running down all circuits. However, as in prior work, we find major differences in similarity depending on whether we measure overlap and cross-task faithfulness (Hanna, Pezzelle, and Belinkov, 2024). Measuring overlap shows low formal-functional similarity, but also low formal-formal similarity; in contrast, cross-task faithfulness suggests clear formal-functional dissimilarity and formal-formal similarity. Overall, we interpret this as showing the existence of a loose collection of formal mechanisms, where circuits drawn from this set are more able to solve formal tasks than circuits drawn from outside it. Not all of these mechanisms are strictly necessary for all formal tasks; a single formal task may have multiple solutions within this flexible collection. Future work could study more tasks, in order to precisely define the limits of LLMs’ formal mechanisms, and further the use of mechanistic interpretability to understand the relationship between LLMs and the brain. Future work could also study the emergence of potentially shared mechanisms during LLM pretraining.
@@ 10 Limitations
In this study, we analyzed five formal and five functional tasks, limited by the difficulty of crafting tasks that fit into the circuits framework; studying more tasks would help confirm that formal and functional tasks do indeed have distinct patterns in cross-task faithfulness. Moreover, we have studied circuits for these tasks found at two faithfulness thresholds, 85% and 90%; varying these thresholds further could still cause our results to change.
@@ Acknowledgments
The authors would like to thank Kyle Mahowald and Anna Ivanova for helpful discussions about this paper, and Raquel Fernández for reading it and providing insightful feedback. The authors also thank the Dialogue Modeling Group at the University of Amsterdam, and the Technion NLP group for their many useful comments and suggestions. We thank anonymous reviewers for directing us to the neuroscientific literature on degeneracy and NPIs.
MH is supported in part by an OpenAI Superalignment Grant. This work was conducted in part during a research visit funded by the ELIAS Mobility Program (Grant Agreement No. 101120237). This research was supported by an Azrieli Foundation Early Career Faculty Fellowship and by Open Philanthropy. This research was partly funded by the European Union (ERC, Control-LM, 101165402). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@
However, for some tasks (Gendered Pronoun, SVA, NPI, and Greater-Than), there is no one correct answer, and that answer need not always be the top output. Consider the case of SVA with the input The keys on the cabinet…; while are is clearly more correct than is, or any other singular-conjugated verbs, much of the model’s probability may be distributed to words like that. Such words are neither right nor wrong, so we should not penalize models for outputting them. In such cases, we score each example as correct if models assign more probability to the right token (or set thereof) than to the wrong token (or set thereof); alternatively, we could use accuracy with respect to the first token that is either correct or incorrect, rather than neither. Results for the tasks shared across all models are in Table˜3. Accuracies tend high, except for IOI and Entity Tracking.
@@
Estimating Overlap with Dummy Circuits
The hypergeometric modeling approach is mathematically interesting, but a poor model of circuits. Circuits do not consist of randomly chosen edges: rather, they consist of paths from the inputs of the graph to the logits. Moreover, each node in a circuit tends to have a limited number of important, high contribution edges; that is, the distribution of edge importance is not uniform. How can we simulate the random selection of such circuits?
We do so using the following procedure. For each task, with a corresponding circuit, we construct a dummy circuit. This consists of a computation graph where each edge has a score drawn from a log normal distribution; we find that this simulates well the tendency of nodes to have just a few contributing edges. We then take the top-$n$ edges from the dummy circuit by (absolute) score, where $n$ is the number of edges used when applying top-$n$ search to find the corresponding task circuit. We prune both the task and dummy circuit, ensuring that neither has any nodes that lack outgoing or incoming edges, and that both sides of each edge are in the circuit.
We construct 20 dummy circuits for each task and model, and compute the average IoU between the task and dummy circuit. We find that the task and dummy circuits have rather low IoUs: most fall below 0.02, less than a fifth of our median IoU of 0.11 between pairs of real circuits. So, even with a more realistic model of random circuits, the overlap we find lies notably above the mean.
@@
More Purely-Functional Tasks We also study two additional tasks that are purely functional. The first is Nikankin et al.’s (2024) basic Math task, which consists of simple arithmetic problems involving addition, subtraction, and multiplication, such as 25 + 3 =. Corrupted examples share the same operator, but different operands. We measure performance via the difference in logit assigned to the clean and corrupted equations’ answers. Presented in purely symbolic form, this task should engage purely functional (formal reasoning) areas of the model, if such exist. This task is highly sensitive to tokenization, [^5] LLMs differ widely in how they tokenize numbers; while Llama-3 tokenizes numbers up to 3 digits in length as one token, other models (e.g. Gemma-2) tokenize them digit-by-digit, and others tokenize them in groups of 2. Single-token prediction tasks are most compatible with circuit analysis, so we use Llama-3. and cannot easily be ported to new models, so we study it in Llama-3 (8B) alone.
The second is a string manipulation task called Echo, inspired by a task of the same name from Hupkes et al. (2021). In this task, models see 4 distinct tokens randomly sampled from the model’s tokenizer, and must repeat the last token; a successfully solved example looks like $t_{1}$ $t_{2}$ $t_{3}$ $t_{4}$ : $t_{4}$”. We give the model two solved examples (2-shot), and then have it finish an incomplete example. In corrupted examples, the final token ($t_{4}$) is replaced with another, distinct token. We measure performance via the logit assigned to the clean and corrupted examples’ final token. Since the tokens are randomly sampled, there should be no trace of linguistic structure; this task is most akin to a pattern-recognition formal reasoning task.
We find the circuits for these two tasks, and compare them to our existing task circuits using IoU, as done previously. Our results on these two tasks (Figure˜11) are similar to those of other functional tasks. Neither Math nor Echo is particularly similar to any other task. Notably, although one might hypothesize that they are supported by the same mechanisms, Math and Greater-Than do not have especially high overlap. Echo also overlaps slightly more with Colored Objects, IOI, Entity Tracking, and Hypernymy than with other tasks. In general, IoU between Math / Echo and formal tasks is low but non-zero (0.09 on average). We take this as evidence that the overlap between formal and functional linguistic tasks is not due to the latter containing the former.
[…]
Circuit Overlap Between Similar Tasks How much do circuits for similar tasks overlap? In the main text, we study Capital-Country and Country-Capital, finding a high IoU (0.5) and near-perfect cross-task faithfulness. However, though these tasks seem , they need have maximally similar metric; past work suggests that models trained on “A is B” relations do not necessarily learn “B is A”, indicating that these two tasks could rely on different mechanisms (Berglund et al., 2024). Could more similar tasks have more similar circuits? And what is the range of IoU and cross-task faithfulness on non-identical circuits? In the following two sections, we compute the similarity of even more similar tasks. In general, we find that very high IoUs ($\geq 0.5$) are uncommon, while very high cross-task faithfulness values are not. This suggests that even relatively low levels of edge IoU may indicate a higher degree of similarity than one might intuitively expect.
Greater-Than Variants
In past work, Hanna, Liu, and Variengien (2023) found high cross-task faithfulness between the Greater-Than task and variants that phrased the task as a statement about prices (Greater-Than-Price: “The price of the purse ranges from \$1842 to \$18…”) or as a sequence of increasing numbers (Greater-Than-Sequence: “1734, 1745, 1788, 1801, 1842, 18…”). We construct the same tasks for Llama-3 (8B), and compute the edge IoU between these tasks. Surprisingly, the edge IoU ranged from 0.25 - 0.31, lower than expected for such similar tasks with high reported cross-task faithfulness.
Functional vs. Purely-Functional Tasks
In Section˜6.2, we studied purely-functional versions of functional tasks, which contain no natural language structure. We can also study the overlap between similar tasks by comparing the circuits for functional tasks with those for their purely-functional counterparts. We do this, and compute (edge) IoU and cross-task faithfulness.
Figure 12: Edge IoU (left) and cross-task faithfulness (right) between pairs of functional tasks and their purely-functional (PF) counterparts. In all cases, there is notable similarity between each functional and purely-functional pair. However, the similarity between normal functional tasks and their purely-functional counterparts is much stronger when measured using cross-task faithfulness.
Our results (Figure˜12) suggest that there is indeed overlap between functional tasks and their purely-functional counterparts. In the case of Edge IoU, this IoU is modest but clearly present: IOI, Colored Objects, and Entity Tracking overlap with their purely-functional counterparts. The same is true for Country-Capital and Capital-Country. However, the maximum IoU between a functional and purely-functional task is around 0.5. In the case of cross-task faithfulness, these trends emerge even more strongly: functional tasks have cross-task faithfulness of near 1.0 on their purely-functional counterparts. The major exception to this is IOI (PF) and the original IOI; notably, the former is a much simplified version of the latter. There are discrepancies between the two metrics: IoU indicates very low similarity between Greater-Than and its purely-functional counterpart, while cross-task faithfulness suggests that their circuits are highly similar. This highlights the differences in the two metrics. The overall difference in ranges also suggests that we should consider even small levels of IoU to be significant, given the compressed range of values it takes on.
[…]
Replication with Higher Faithfulness Threshold
In order to ensure that our results are result to the chosen 85% faithfulness threshold, we ran our experiments again, with a higher, 90% faithfulness threshold. Notably, the results are the same, in both the heatmap (Figure˜13) and dendrogram (Figure˜14) analyses at the edge level, as well as at the node / neuron level (not shown).
Figure 13: IoU and cross-task faithfulness heatmaps using a 90% threshold. The same trends emerge as in the original 85% analysis. There is low formal-functional similarity in the edge IoU case, but also low formal-formal similarity. In the cross-task faithfulness case, there is a stronger formal-formal overlap.
Figure 14: IoU and cross-task faithfulness dendrograms using a 90% threshold. The same trends emerge as in the original 85% analysis. The edge IoU dendrogram yields imperfect groupings (with Hypernymy and Greater-Than swapped), while the cross-task faithfulness dendrogram recovers the formal-functional distinction perfectly.
[…]